if a,b,c belongs to R and a²+5/2b²+c² =2ab + bc + ca ,then find a + 2b + c.

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
As given,
a²+5/2b²+c² =2ab + bc + ca
a^2 + c^2 -2ac + a^2 + 4b^2 – 4ab + c^2+ b^2-2bc = 0
=>(a-c)^2 + (a-2b)^2+ (b-c)^2 = 0
a=c , a=2b, b=c 
a= b = c = 0
a+ 2b + c = 0
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)